How do you work
{y=2x
{y=-6x+4

Found 3 solutions by Alan3354, stanbon, eperette:
Answer by Alan3354(69443)

How do you work
{y=2x
{y=-6x+4
----------
There are several methods, but since you have 2 that say =y, just set them equal to each other.
2x = 6x+4
-4x = 4
x = -1
----------
y = 2x
y = -2
Always take the easiest route.

Answer by stanbon(75887)

How do you work
{y=2x
{y=-6x+4
-----------
2x = -6x+4
8x = 4
x = 1/2
-----------
Substitute to solve for "y":
y = 2(1/2) = 1
=========================
Cheers,
Stan H.

Answer by eperette(173)

How do you work
{y=2x
{y=-6x+4
you can do this by substitution since you already know y=2x, you can replace the y in the second equation by 2x...
y = -6x + 4
2x= -6x + 4
2x+ 6x = -6x + 6x +4
8x = 0 + 4
8x = 4
8x/8 = 4/8
x = 1/2
now that you know x go back and find y by substitution also...
y = 2x
y = 2(1/2)
y = 2/2
y = 1
so your answer is....
(1/2, 1)